For a data set of the pulse rates for a sample of adult females, the lowest pulse rate is 36 beats per minute


1
Expert's answer
2021-11-09T00:06:11-0500

As the question is incomplete, we make it as follows:

For a data set of the pulse rates for a sample of adult females, the lowest pulse rate is 36 beats per minute and the highest rate is 80 beats per minutes. Find the mean.

Solution:

Mean "=\dfrac{36+80}2=\dfrac{116}2=58" beats per minute.